Parth (parthkohli):

Let \(a\) be a positive real number. If the following equation has a real root... \(x^6 + 3a x^5 + (3a^2 + 3)x^4 + 6ax^3 + (3a^2 + 3)x^2 + 3ax + 1 + (ax)^3 = 0\) Then find the minimum possible value of \(a\).

Parth (parthkohli):

@Callisto There.

Parth (parthkohli):

@perl

Parth (parthkohli):

Hint: there's a pattern hidden in there. :)

ganeshie8 (ganeshie8):

any other usefufl hint ?

Parth (parthkohli):

Do you see the pattern? You should be able to solve it if you do...

ganeshie8 (ganeshie8):

hmm i can only think of vieta formula and it looks hard

Parth (parthkohli):

There's a thing with the coefficients. They repeat in a fascinating manner.

OpenStudy (anonymous):

(x-a)^6=0

OpenStudy (anonymous):

x=a

Parth (parthkohli):

@Rishi123 I don't see how you reduced it to \((x-a)^6 =0\)

Parth (parthkohli):

Let \(a\) be a positive real number.

OpenStudy (anonymous):

Are you supposed to use grouping to solve for the polynomial?

Parth (parthkohli):

Yes.

OpenStudy (perl):

you can rewrite the lefthand side as ( x^2 + ax + 1)^3 = 0

Parth (parthkohli):

You're damn right.

OpenStudy (perl):

then use derivative to minimize that, i think

Parth (parthkohli):

No need.

ganeshie8 (ganeshie8):

is that the pattern ? i wouldn't have figured that without wolfram :O

OpenStudy (perl):

you can look at the discriminant ?

OpenStudy (perl):

that has a real solution if a^2 -4 (1)(1) > 0

OpenStudy (perl):

>=

Parth (parthkohli):

Pretty sure that perl has looked at Wolfram too, but my method was different.

OpenStudy (jtvatsim):

Applied mathematicians unite! What sort of pure math madness is this? jk :)

OpenStudy (perl):

so the minimum a, is 2?

Parth (parthkohli):

God, I just lost all of my work...

Parth (parthkohli):

Step 1: divide both sides by \(x^3\). You do this when you have mirroring coefficients.

Parth (parthkohli):

\[\left(x^3 + \dfrac{1}{x^3}\right) + 3a \left( x^2 + \dfrac{1}{x^2}\right) + (3a^2 + 3) \left(x + \dfrac{1}{x}\right) + 6a + a^3 = 0\]

Parth (parthkohli):

Let \(t = x + \dfrac{1}{x}\). Then \(x^2 + \dfrac{1}{x^2} = t^2 - 2\). And \(x^3 + \dfrac{1}{x^3}= t^3 - 3t\)

Parth (parthkohli):

Plug those in and then with the cancellations and stuff, you'd be left with\[t^3 + 3at^2 + 3a^2 t + a^3 = 0\]\[\Rightarrow (t + a)^3 = 0\]

ganeshie8 (ganeshie8):

wow! nice thats pretty patterny xD

OpenStudy (perl):

what was the minimum value for a?

ganeshie8 (ganeshie8):

\(x + \frac{1}{x} = -a \implies x^2 +ax+1 = 0 \) \(D \ge 0\) yields \(a^2 - 4 \ge 0\) as desired

OpenStudy (perl):

yes thats what I got too :)